OpenStudy (anonymous):

A floor has two square-shaped designs. The area of the second square-shaped design is nine times greater than the area of the first square-shaped design. Which statement gives the correct relationship between the lengths of the sides of the two squares?

OpenStudy (anonymous):

these are the options... The length of the side of the second square is 3 times greater than the length of the side of the first square. The length of the side of the second square is 12 times greater than the length of the side of the first square. The length of the side of the second square is 9 times greater than the length of the side of the first square. The length of the side of the second square is 6 times greater than the length of the side of the first square.

OpenStudy (anonymous):

since \(A=l^2\) and \(A'=l'^2\) when \(A=9A'\) sub them in to get the relationship between l and l'

OpenStudy (anonymous):

what am i supposed to sub in?

OpenStudy (anonymous):

the l and l ' like this: \(A=9A'\) \(l^2 = 9 l'^2\)

OpenStudy (anonymous):

im like really slow and have no idea wut im doing.-_-

OpenStudy (anonymous):

lol. i guess i'll try to do it step by step. since \(A=l^2\), \(A_1=l_1^2\) (this is the first square)............(1) \(A_2=l_2^2\) (this is the second square)........(2) since the area of the second square is nine times the first, (it's the only given condition, so we start from that.) \(A_2=9A_1\) subbing (1) and (2) into it, we get, \(l_2^2=9l_1^2\)| what did you get after taking thesquare roots of both sides?

OpenStudy (anonymous):

sorry im back and thank you.

OpenStudy (anonymous):

Just take the square root both the sides and tell us what you got as @Shadowys said above..

OpenStudy (anonymous):

Getting ?? @Brianna9898

OpenStudy (anonymous):

I have no idea!

OpenStudy (anonymous):

Are you getting till here" \[l_2^2=9 l_1^2\]

OpenStudy (anonymous):

why is there a 1 & 2 at the bottom

OpenStudy (anonymous):

See when you will take square root you will get like: \[\large \sqrt{l_2^2} = \sqrt{9 l_1^2}\] Can you tell what is this: \[\large \sqrt{l_2^2} = ??\]

OpenStudy (anonymous):

... and how am I supposed to square root an l ??

OpenStudy (anonymous):

Oh that.. \(l_1\) is showing length of first square. \(l_2\) is for length of second square.

OpenStudy (anonymous):

you can actually.. See, What is square root of this: \[\large \sqrt{2^2} = ??\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Similar way what will be square root of this: \[\large \sqrt{l_2^2} = ??\]

OpenStudy (anonymous):

\[l\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Or you can say \(l_2\).. Good..

OpenStudy (anonymous):

Similarly can you tell for: \[\large \sqrt{9l_1^2} = ??\]

OpenStudy (anonymous):

so you have to keep the 2 at the bottom

OpenStudy (anonymous):

the one and two are sub scripts to differentiate between the first length and the second, thus, 1 and 2 respectively.

OpenStudy (anonymous):

see, 1 and 2 is differentiating lengths of the two square you are given with, so don't think here of just l think here of \(l_1\) and \(l_2\)..

OpenStudy (anonymous):

so it would be \[9l _{1} ??\]

OpenStudy (anonymous):

You forgot to take square root of 9. \(l_1\0 is good though..

OpenStudy (anonymous):

What is square root of 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yep after square root you will get like: \[\large l_2 = 3 \times l_1\]

OpenStudy (anonymous):

So, which answer choice is this?

OpenStudy (anonymous):

the first one?

OpenStudy (anonymous):

Well Done...

OpenStudy (anonymous):

And give all the thanks to @Shadowys

OpenStudy (anonymous):

yayyy! thank you guys for helping me!

OpenStudy (anonymous):

I don't even know how to give a medal on this thing

OpenStudy (anonymous):

Are you seeing best response after shadows post ?/ Just click that..

OpenStudy (anonymous):

oh okay, can I give a medal to two ppl?

OpenStudy (anonymous):

No, just only one..

OpenStudy (anonymous):

On one post you can give medal to one only..

OpenStudy (anonymous):

thanx for the medal @Shadowys !